### v2

# AXIS: Efficient Human-Agent-Computer Interaction with API-First LLM-Based Agents

Abstract
Multimodal large language models (MLLMs) have enabled LLM-based agents to directly interact with application user interfaces (UIs), enhancing agents’ performance in complex tasks. However, these agents often suffer from high latency and low reliability due to the extensive sequential UI interactions. To address this issue, we propose AXIS, a novel LLM-based agents framework that prioritize actions through application programming interfaces (APIs) over UI actions. This framework also facilitates the creation and expansion of APIs through automated exploration of applications. Our experiments on Microsoft Word demonstrate that AXIS reduces task completion time by 65%-70% and cognitive workload by 38%-53%, while maintaining accuracy of 97%-98% compared to humans. Our work contributes to a new human-agent-computer interaction (HACI) framework and explores a fresh UI design principle for application providers to turn applications into agents in the era of LLMs, paving the way towards an agent-centric operating system (Agent OS).
AXIS: Efficient Human-Agent-Computer Interaction with
API-First LLM-Based Agents
Junting Lu1*, Zhiyang Zhang2††thanks: Equal Contribution. Work is done during the internship at Microsoft., Fangkai Yang3††thanks: Corresponding author., Jue Zhang3, Lu Wang3, Chao Du3, Qingwei Lin3, Saravan Rajmohan3, Dongmei Zhang3, Qi Zhang3 1Peking University, 2Nanjing University, 3Microsoft aidan.lew.37@stu.pku.edu.cn
1 Introduction
As personal computers and mobile devices become indispensable in daily life, application industries face pressure to rapidly evolve software with new features to meet growing demands Ruparelia (2010); Abrahamsson et al. (2017). However, to use a new application effectively, users must first spend time familiarizing themselves with the user interface (UI) and its functionalities, increasing users’ cognitive burden Van Merrienboer and Sweller (2005); Biswas et al. (2005); Plass et al. (2010); Darejeh et al. (2022). Large language models (LLMs) Ouyang et al. (2022); Achiam et al. (2023); Dubey et al. (2024) have demonstrated near-human capabilities in reasoning, planning, and collaboration and are highly promising in completing complex tasks Huang and Chang (2022); Wei et al. (2022); Mandi et al. (2024). Since then, researchers have been leveraging multimodal large language models (MLLMs) Yin et al. (2023); Durante et al. (2024); Zhang et al. (2024b) to operate software applications with vision capability Wu et al. (2023); Zheng et al. (2024). Recent works Zhang et al. (2023a); Wang et al. (2024a); Zhang et al. (2024a); Zheng et al. (2024) utilize MLLMs to design LLM-based UI agents capable of serving as user delegates, translating user requests expressed in natural language, and directly interacting with the application’s UIs to fulfill users’ needs without a deep understanding of UIs and functionalities.
However, just like the transition from steam-powered to electric-powered industry took much more than replacing central steam engines with electric motors in the factories, simply building LLM-based UI agent cannot magically deliver a satisfying and worry-free user experience. In particular, today’s application UIs are designed for human-computer interaction (HCI) Lewis (1998); Bradshaw et al. (2017), which often involves multiple UI interactions for completing a single task. For instance, inserting a 22 table in an Microsoft Word document requires a sequence of UI interactions: “Insert Table 22 Table”. Although the HCI-based design suits the habits of humans, training LLM-based UI agents to emulate such interactions would pose quite a few challenges that are difficult to overcome.
The first challenge for the LLM-based UI agent is high latency and long response time. Each individual UI interaction step requires one LLM call to reason which UI to interact with. A task involves multiple UI interaction steps can thus incur considerable time and monetary costs. The LLM call latency is also positively correlated with the number of processed tokens Levy et al. (2024); Wang et al. (2024d); Egiazarian et al. (2024). To ensure the LLM returns high-quality outputs, the LLM-based UI agent must pass a large volume of UI information to precisely describe the current state, which also increases latency in each call.
The second challenge lies in the domain of reliability. Studies have shown that LLMs are prone to hallucinations in generating responses and selecting correct UIs when a long chain of UI interactions is required Bang et al. (2023); Dhuliawala et al. (2023); Zhang et al. (2023b); Guan et al. (2024). During long sequential calls, the chance of selecting a wrong UI control or hallucinating a non-existent UI for interaction increases with each reasoning step Zhang et al. (2024a), resulting in compounding errors and task failure Chen et al. (2024); Zhao et al. (2024). Lastly, the LLM-based UI agent also faces the challenge of UI generalization. While recent works had made advancements in UI grounding Cheng et al. (2024); Rawles et al. (2024b); Bai et al. (2023), how the LLM-based UI agents handle interactions with applications whose UIs are not included in the pretraining stage of LLMs remains a critical obstacle without good solutions.
To address these challenges faced by LLM-based UI agents, we highlight the necessity of application programming interfaces (APIs) within the new human-agent-computer interaction (HACI) paradigm. As an intermediate layer between human and computer, the LLM-based agent should understand user requests in natural language and operate computers/applications by prioritizing APIs over human-like UI interactions. Inspired by this, we propose AXIS: Agent eXploring API for Skill integration, a self-exploration LLM-based framework capable of automatically exploring existing applications, learning insights from the support documents and action trajectories, and constructing new APIs111The new APIs are also referred as “Skills” in Section 4, and we use the term “API” loosely here to differentiate from the UI. based on the existing ones. AXIS helps build API-first LLM-based agents that replace UI agents to prioritize API calls over unnecessary multistep UI interactions in task completion. Regular UI interactions are only called when the related APIs are unavailable. Figure 1 shows a task completion with the UI agent and API-first agent. Compared to the UI agents, API-first agents require fewer tokens and can obtain more accurate, code-formatted responses from LLMs with low latency and high reliability.
Our work makes the following contributions:
We propose a novel HACI paradigm along with an implementation framework, AXIS, which enables API-first LLM-based agents to explore an application’s available APIs and construct new ones as needed. This approach facilitates the straightforward transformation of any application into an autonomous agent by wrapping it with an API set and adopting a simplified UI design. This paradigm not only addresses practical challenges faced by application providers but also paves the way toward a full-fledged Agent Operating System (Agent OS) Zhang et al. (2024c); Mei et al. (2024); Wu et al. (2024).
We mitigate cognitive load and reduce the learning effort required for complex interactions by replacing multi-step UI operations with efficient API calls. Our experiments on Microsoft Word tasks Microsoft365 (2024a) show that AXIS significantly improves task completion rates while alleviating the cognitive burden on users.
We conduct comprehensive performance evaluations and an extensive user study to validate the efficiency, reliability, and practical applicability of AXIS in real-world scenarios.
2 Related Work
2.1 LLM-based UI Agent
LLM-based agents are designed to utilize the advanced context understanding and reasoning skills of LLMs to interact with and manipulate environments with human-like cognitive abilities † et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu et al. (FAIR); Xi et al. (2023); Liu et al. (2023); Wang et al. (2024b). The advent of MLLMs Yin et al. (2023); Durante et al. (2024); Zhang et al. (2024b), including GPT-4o OpenAI (2024) and Gemini Team et al. (2023), expands the research landscape for LLM-based UI agents. LLM-based UI agents have been applied to multiple areas such as mobile platforms Yan et al. (2023); Zhang et al. (2023a); Wang et al. (2024a), Web Song et al. (2024) and OS Wang et al. (2024c); Zhang et al. (2024a); Zheng et al. (2024); Tan et al. (2024); Hong et al. (2024); Cheng et al. (2024). LLM-based UI agents mimic human interactions, but existing UIs are designed for human-computer, not agent-computer, interactions, leading to inefficiencies in repetitive tasks. APIs offer a more efficient alternative by reducing unnecessary UI steps. We explore leveraging APIs for LLM-based agents and propose new UI design principles for the LLM era.
2.2 Agent Operating System
To support the completion of complex tasks with minimal human interventions, emerging works have explored the possibility of developing an agent operating system (Agent OS) fully supported by LLMs Mei et al. (2024); Wu et al. (2024); Zhang et al. (2024c); Xie et al. (2024); Rawles et al. (2024a). In the industry, commercial Agent OSes Apple (2024); Microsoft (2024); Huawei (2024); Honor (2024) are evolving to become more accessible and productive for customers with the potential of leading a new era of HCI. Existing Agent OSes typically divide complex tasks into sub-tasks and assign them to applications. However, LLM-based agents still rely on human-like UI interactions, such as clicking and swiping, which are less efficient than API calls. Additionally, when processing a task, the LLM-based UI takes control away from the user.
2.3 UI design in LLM era
UI design is an essential part of HCI and requires highly specialized expertise along with iterative rounds of feedback and revision Stone et al. (2005). With LLMs, UI design can be further empowered with automated procedures of design, feedback and evaluation. Duan et al. (2024) use LLM-generated feedback to automatically evaluate UI mockups. Similiarly, SimUser Xiang et al. (2024) leverages LLMs to simulate users with different characteristics to generate feedback on usability and provide insights into UI design. MUD Feng et al. (2024) utilize LLMs to mimic human-like exploration to mine UI data from applications and employs noise filtering to improve quality of UI data. Existing UI designs follow the traditional HCI paradigm rather than the HACI paradigm central to Agent OS. Using the AXIS framework, we explore applications, identify essential UIs, and determine which UI components can be replaced by API calls for LLM-based agents.
3 Preliminary
Environment. In the context of AXIS, the environment refers to the collection of interactive entities within the exploration scope of the agents. In our case, these entities primarily consist of applications running on the Windows operating system, with Microsoft Word being our focus. Applications in the environment often share common elements, such as controls Zhang et al. (2024a) and XML elements obtained after unpacking. To facilitate the observation and interaction between agents and the environment, we have designed two general interfaces: state() to return the environment state and step() to execute agent actions, respectively. Details of interfaces are in Appendix A.
Skills. A skill in AXIS is a structured unit designed to accomplish a specific task within the environment. It is a high-level representation of UI- and API-based actions, with priority given to API actions222If the skill can be represented with UI or API actions, the skill is represented in API-only actions.. Following the design of tool usage and function call Cai et al. (2023); Wang et al. (2023), each skill consists of three components: description, skill code, and usage example. Appendix B.1 shows details of each component.
Skill Types and Hierarchy. Following a versatile design principle, the skills in AXIS can be categorized into five types based on the composition of their code fragments: Atomic UI Skill, Atomic API Skill, Composite UI Skill, Composite API Skill, and API-UI Hybrid Skill, the details are shown in Appendix B.2 Table 6. Additionally, we define “skill hierarchy” as the number of skills contained. A single atomic skill thus has a skill hierarchy of 1. Based on their code composition and hierarchy, skills in AXIS can be nested. For example, skill A can call skill B, which in turn calls skill C, forming a hierarchy of depth 3.
4 Design of AXIS
We develop AXIS as a framework that can automatically explore within existing application environments, learn insights from exploration trajectories, and consolidate available insights and learned knowledge into actionable “skills”. Illustrated by Figure 2, the AXIS system is composed of three crucial stages: trajectory collection, skill generation, and skill validation. The trajectory collection stage collects interaction trajectories in task completion, and then the skill generation stage generates skills from these trajectories and translates into skill code. The skill validation stage validates the skill code before it is added in the skill library to reduce hallucination and maintain generalizability.
4.1 Stage I: Trajectory Collection
“Experience is the mother of wisdom”. Following the approach in Wang et al. (2023), we let the agent practice acquiring skills in the application environment. In this stage, we designed two practice modes: Follower Mode and Explorer Mode, which are task-oriented collections with and without task descriptions, respectively. The specific prompts refer to Appendix LABEL:sec:traj-prompts.
Follower Mode. The Follower Agent extracts tasks and instructions from the application help documents. The agent processes structured inputs: task requirements, step-by-step instructions, environmental context, and the current skill library. It strictly adheres to instructions while interacting with seed application files, collecting execution trajectories. Action selection employs the ReAct mechanism Yang et al. (2023), integrating action history, step descriptions, and environmental states. Initially limited to basic actions (see Table 7), the follower agent is able to perform more complex tasks with the expansion of the skill library through the other parallel processing stages.
Explorer Mode. Explorer mode differs from the follower mode by leveraging the brainstorming capability of LLMs to generate various step-by-step instructions. The Explorer agent observes environment states and action history, selecting actions from the skill library without following predefined task guides. To enhance state diversity, we implemented:
Initial State Diversity: Random replay of Microsoft Word files up to intermediate steps, providing varied starting points.
Vertical/Horizontal Exploration: A random “dive into” strategy balancing continuity (vertical) and breadth (horizontal) in functionality exploration. “Vertical” means the agent will select UI controls at lower levels of the current control tree for its next action, while “Horizontal” means the agent will tend to select UI controls at the same level or switch to other menus.
Skill Proficiency Levels: Three explorer levels (low/medium/high) guided by Microsoft Office Specialist (MOS) certification curriculum333https://learn.microsoft.com/en-us/credentials/certifications/mos-word-2019/, with system prompts tailored to their app familiarity.
4.2 Stage II: Skill Generation
This stage runs in parallel with Stage I and is responsible for converting the trajectories into structured skills. This process involves three LLM-based agents. The specific prompts for these agents are detailed in Appendix LABEL:sec:skillgen-prompts.
Monitor. The collected trajectories often contain complex and multiple interaction steps, making it challenging to convert them into a single skill without enough basic skills available in the skill library. The monitor examines the skill library, extracts meaningful trajectory segments, and consolidates them into natural language-based skill insights.
Generator. The Generator agent faithfully combines the skill insights with their corresponding trajectory segments and generates executable skill code. As shown in Figure 2, the skill code mainly consists of stacked actions from the trajectories, where the original action parameters are converted into placeholders. Along with the skill code generation, it also produces corresponding skill usage examples and descriptions, with examples shown in Table LABEL:tab:skill-trans. Note that since the skill code is a faithful reproduction of the trajectory, which may contain numerous UI operations, the skill code essentially becomes a combination of UI actions, leaving room for speed optimization.
Translator. As the skill code from the generator contains numerous UI operations, a “UI” to “API” translation is needed. The Translator agent accomplishes this task by connecting to the retrieval-augmented generation (RAG) module Gao et al. (2023), consulting the application’s documentation and the current skill library to convert UI operation code segments to APIs. As shown in Figure 2, it converts UI operations into direct modifications of target content within the application. Table LABEL:tab:skill-trans shows a translation example.
4.3 Stage III: Skill Validation
Due to potential LLM hallucinations and cumulative errors in multi-agent transmission, generated skills may have syntax or functionality issues. To address this, we implement static and dynamic validation, with dynamic validation following static validation.
Static Validation. Static validation utilizes structural checks to verify the compatibility of the skill code with the skill executor. It examines whether the skill’s parameters contain the mandatory parameters (such as the executor instance and args list), whether the methods and properties of the executor are correctly invoked in the code, and whether any non-existent skills are imported when reusing the skill. Skills that do not meet these formal requirements are returned for revision.
Dynamic Validation. Dynamic validation evaluates the performance of a skill in a real environment. It consists of a validator and an evaluator. The validator generates various input parameters to test the skill’s generalizability, and the evaluator checks whether the test is successfully completed by examining the final state. The specific prompts are detailed in Appendix LABEL:sec:skillval-prompts, with detailed description of the validator and evaluator in Appendix LABEL:appendix:dynamicvalidator.
5 Feasibility Study
To validate the usability and effectiveness of the AXIS framework, we conduct a feasibility study. We first use AXIS to explore Microsoft Word and discover 73 skills. Then we extract 50 tasks from the wikihow 444https://www.wikihow.com/Use-Microsoft-Word page “Use Microsoft Word” and the official Microsoft Word website 555https://support.microsoft.com/en-us/word. These tasks were executed using both AXIS and UI Agent, and the results were analyzed and compared. AXIS also enable turn an application into an agent by simply wrapping the application with an API set and adopting a simpler UI design suitable for HACI, which is presented in Appendix LABEL:sec:extensive.
5.1 SKill Exploration
Before the exploration, AXIS is provided with the initial skill library which composed of 6 basic actions as shown in Table 7. Then, 347 seed files are used for AXIS to explore. After the exploration, AXIS discovers 73 skills with different hierarchies. Majority of the skills (44) discovered have a skill hierarchy 1. The rest is composed of 24 skills with hierarchy 2, 3 skills with hierarchy 3, and 2 skills with hierarchy 4. Table LABEL:tab:explored displays several successfully validated skills discovered during the exploration process.
5.2 Task Completion
We evaluate UI Agent (represented by UFO Zhang et al. (2024a) for its superior Word task performance) and AXIS on 50 Word-related tasks using explored skills. Table 1 presents comparative results, including completion time, success rates, step counts, and LLM backend costs (GPT-4o, version 20240513) for both agents.
In terms of execution time, AXIS demonstrates superior performance, completing tasks in 29.9 seconds on average - twice as fast as UI Agent’s 59.5 seconds. It also achieves higher success rates and, through its skill-based abstraction, requires fewer steps per task, resulting in lower costs compared to the UI Agent UFO.
| Metric | UI Agent | AXIS | P - Significance |
| Time(s) | 59.5 | 29.9 | u>a (p < 0.001) |
| Success Rate(%) | 52.0 | 84.0 | u<a (p < 0.001) |
| Steps | 3.2 | 2.0 | u>a (p < 0.01) |
| Cost($) | 0.4 | 0.2 | u>a (p < 0.001) |
To investigate AXIS’s efficiency, we analyze action types and API usage patterns. As shown in Table 2, AXIS employs significantly fewer UI actions compared to the UI Agent, while utilizing more API and Advanced API calls (skills with hierarchy level 2). Analysis reveals AXIS predominantly employs integrated API skills for task completion, yielding more API actions (55.7% usage rate, including 23.1% advanced APIs) compared to UI Agent’s 8.1%. This demonstrates AXIS’s API-first approach, leveraging available skills for efficient task execution through skill-action integration.
| Metric | UI Agent | AXIS |
| Total UI actions | 103 | 48 |
| Total API actions | 9 | 39 |
| API usage rate(%) | 8.1 | 55.7 |
| Advanced API usage rate(%) | - | 23.1 |
6 User Study
We conduct an extensive user experiment to evaluate the performance of AXIS. The experiment and evaluation metrics are designed to explore the following research questions (RQs) on the role of LLM-based agents in work and daily life scenarios:
RQ1: Does the LLM-based agent lower the cognitive load of the users and make them have less effort to learn?
RQ2: Does the LLM-based agent enhance the efficiency of users?
RQ3: What are the differences between a UI Agent and an API-based Agent in user experience?
In our user experiment, participants are asked to complete specified tasks within an application through three methods: manually, with the assistance of a UI Agent, and with the assistance of AXIS. The entire process is recorded. Microsoft Word is chosen as the experimental application considering its popularity in our daily work and life as well as the rich API documentations Microsoft365 (2024b)). Thus five tasks of Word are sampled from both official Word documentation and GPT-generated results, divided into two difficulty levels: low difficulty (L1) and high difficulty (L2). Motivated by the RQs, we set three objectives for the user experiment: (1) To evaluate the cognitive load on participants when completing tasks using different methods. (2) To compare the efficiency and reliability of task completion across the three methods. (3) To assess user preferences regarding the use of different Agents. This study is approved by the Institutional Review Board (IRB) of University.
The details of experiment procedure and participants recruitment are shown in Appendix LABEL:sec:user-study.
| Metric | Task Level | Manual | UI Agent | AXIS | Pairwise Significance |
| Time(s) | L1 L2 | 61.8 167.6 | 104.6 155.5 | 18.2 57.1 | L1: m<u (p < 0.001) L1, L2: a<m (p < 0.001) L1, L2: a<u (p < 0.001) |
| Success Rate(%) | L1 L2 | 100.0 97.5 | 75.0 45.0 | 98.3 95.0 | L1, L2: m>u (p < 0.001) L1, L2: a>u (p < 0.001) |
| Metric | Task Level | UI Agent | AXIS | Pairwise Significance |
| steps | L1 L2 | 6.4 11.1 | 1.0 4.2 | L1: a<u (p < 0.001) L2: a<u (p < 0.001) |
| cost($) | L1 L2 | 0.6 0.9 | 0.07 0.3 | L1: a<u (p < 0.001) L2: a<u (p < 0.001) |
6.1 Experimental Metrics
We collect both subjective and objective metrics in our experiments to evaluate the performance and user experience of different methods.
Subjective Metrics. As detailed in Appendix LABEL:sec:user-study, we conduct four post-task questionnaires (manual and Agent-assisted L1/L2 tasks) using NASA-TLX Hart (1988) metrics: Mental, Physical, and Temporal Demand; Performance; Frustration; and Effort, supplemented by a learning effort metric. Lower scores indicate reduced cognitive load, improved success perception, and decreased frustration/effort. For Agent-assisted tasks (Questionnaires 3-4), additional metrics include Agent fluency/reliability, UI dependency, decision consistency (Agent-user alignment), and perceived completion speed.
Objective Metrics. For objective metrics, we maintain comprehensive experimental logs, including screen recordings of manual and Agent-assisted tasks, decision-making processes, UI interaction paths, task completion time and success rates, LLM backend costs (GPT-4, version 20240513) for Agent operations, and UI dependency measurements for both the UI Agent and AXIS.
| Metric | Task Level | Manual | Agents | Pairwise Significance |
| Mental Demand (0-100) | L1 L2 | 21.3 70.0 | 2.5 7.5 | L1: m>a (p < 0.001) L2: m>a (p < 0.001) |
| Physical Demand (0-100) | L1 L2 | 31.3 57.5 | 5.0 6.3 | L1: m>a (p < 0.001) L2: m>a (p < 0.001) |
| Temporal Demand (0-100) | L1 L2 | 52.5 37.5 | 28.8 35.0 | L1: m>a (p < 0.05) L2: - |
| Performance (0-100) | L1 L2 | 21.2 47.5 | 21.2 26.2 | L1: - L2: m>a (p < 0.05) |
| Frustration Level (0-100) | L1 L2 | 31.3 62.5 | 7.5 10.0 | L1: m>a (p < 0.001) L2: m>a (p < 0.001) |
| Completion Effort (0-100) | L1 L2 | 12.5 35.0 | 17.5 13.8 | L1: - L2: m>a (p < 0.01) |
6.2 Results
Our analysis explores three key aspects: (1) cognitive load reduction, (2) task efficiency of agents, and (3) user preferences between UI agents and AXIS.
Cognitive Load. To assess cognitive load reduction by LLM-based Agents, we analyze NASA-TLX and learning effort scores (Table 5, Figure 3). Results show higher scores for L2 tasks across dimensions, validating our task difficulty classification. Agent-based methods significantly outperform manual approaches in reducing mental/physical demand and frustration, particularly for complex L2 tasks (p<0.05). While L1 task performance differences were minimal, agents notably enhanced users’ success perception in L2 tasks. Analysis reveals consistent user experiences across task complexities (Figure 3 (b)) and significant learning effort reduction, especially for difficult tasks (Figure 3 (c)). These findings answer RQ1: LLM-based agents effectively reduce cognitive load and learning effort, particularly for complex tasks.
Efficiency and Reliability. Our evaluation compares manual, UI Agent, and AXIS approaches using completion time, success rate, steps, and costs (summarized in Tables 3 and 4). AXIS demonstrates superior time efficiency, significantly outperforming both methods (p < 0.001), particularly in complex L2 tasks. While manual methods achieve highest accuracy, AXIS maintain near-human performance, contrasting with UI Agent’s lower accuracy due to UI element positioning and visibility issues. Notably, UI Agents require substantially more steps for deeply nested L2 tasks, whereas AXIS’s API-driven approach enable streamlined execution with fewer steps and lower costs. This answers RQ2: while UI Agents offer modest efficiency gains for complex tasks with reliability challenges, AXIS consistently improves human efficiency with superior reliability.
Affinity Preference. To compare user experiences between UI agents and AXIS, we conduct a subjective evaluation across five aspects (Figure 4). Participants consistently preferr AXIS for its perceived speed, fluency, and reliability across both L1 and L2 tasks. While AXIS’s highly encapsulated API led to less human-like decision-making in simple tasks (L1), its decisions aligned better with human reasoning in complex tasks (L2). AXIS notably improves user’s experienct by reducing UI dependency compared to UI agents’ frequent interface interactions. These findings answer RQ3: AXIS provides superior efficiency, smoothness, and reliability compared to UI agents, with increasing user preference as task complexity grows. User feedback highlights AXIS’s API-first approach as advantageous over UI agents’ intrusive behaviors, offering better control and experience.
7 Conclusion
We present AXIS, a framework that enhances human-agent-computer interaction (HACI) by prioritizing API calls over UI interactions to reduce inefficiencies and cognitive burdens in complex tasks with multimodal large language models (MLLMs). Experiments with Microsoft Word show that AXIS cuts task completion time by 65%-70% and cognitive workload by 38%-53%, while maintaining human-level accuracy. These results demonstrate the potential of API-first LLM-based agents to streamline interactions, reduce latency, and improve reliability, offering a novel approach to faster, more efficient task execution.
8 Ethical Statement
All used datasets collected internally or obtained from external sources, ensuring no infringement on individual or organizational rights. User study participants volunteered and were compensated.
9 Limitations
While AXIS effectively mines APIs and enables efficient human-agent-computer interaction (HACI), further optimization is needed to achieve an Agent OS. Currently, AXIS primarily relies on Python-based APIs, limiting support for applications without native Python interfaces. Additionally, its exploration process requires improvements in stability and efficiency. Future work should focus on developing unified action interfaces to extend HACI to more applications and operating systems while enhancing the framework’s performance and efficiency.
Appendix A Interfaces in Environment
Below are the details of two interfaces: state() and step() to return the environment state and execute agent actions, respectively.
state(): Returns the state of the environment, including detailed information on the current elements within the environment. The environment state encompasses key UI information such as control positions, control types, and selection status, which is consistent with the definition in Zhang et al. (2024a). For applications that can be unpacked, the unpacked XML content is also included as part of the state.
step(): Incorporates a skill executor that allows agents to perform operations within the environment by executing skills. Upon completion, this interface returns the results of these operations.
Appendix B Skills
B.1 Skill Component
Each skill consists of three main components:
Skill Code: A piece of code structured to be compatible with the skill executor. Skill code includes a uniform set of parameters and adheres to the standard PEP 257 documentation. The initial set of skills is generated by restructuring the fundamental APIs from the application provider, and based on these initial skills, AXIS can explore and develop additional new skills.
Description: A description of the skill’s functionality, which assists the LLM in selecting and invoking appropriate skills during task execution.
Usage Example: One or more code examples including typical parameters associated with the code and description. These examples help the LLM in correctly formatting parameters when invoking the skill.
Skill Executor. As discussed above, our application environment incorporates a step() interface to facilitate the interaction between agents and the environment. This interface also hosts the skill executor responsible for executing the skill generated or selected by the agents. The skill executor keeps caching of application documents and simultaneously supports multiple functionalities including locating application controls, invoking methods on those controls, and calling application APIs (independent of controls), to enable the UI actions and API actions in the same time and serve as an efficient foundation for skill-driven operations.
B.2 Skill Type
The skills in AXIS can be categorized into five types based on the types of the code fragments: Atomic UI Skill, Atomic API Skill, Composite UI Skill, Composite API Skill, and API-UI Hybrid Skill. There are some examples in Table 6.
| Type | Description | Example | Feature coverage |
| Atomic UI skill | Composed of one basic UI action. As the most primitive skills, Atomic UI skills are stacked and transformed during the exploration process to form new skills. | click_input | click on different UI controls |
| Atomic API skill | Composed of one basic API actions. Unlike UI actions that depend on UI controls for execution, API actions can be executed without the need of interacting with any UI elements. | select_text | select text content in the canvas. |
| Composite UI skill | Composed of multiple atomic UI actions or composite UI actions. Composite UI skill are formed by a simple stacking and combination of UI actions. | search_for_help | clicking the search box and then editing text. |
| Composite API skill | Composed of multiple atomic API actions or composite API actions. This type of skills often represents a higher-level combination of functions. | insert_header_footer | insert header and footer with specified contents by API, which is equal to sequential UI actions "Insert->Header->footer edit->Footer->footer edit". |
| API-UI hybrid skill | Composed of both API actions and UI actions. API-UI hybrid skills sometimes appear as intermediate states during skill exploration and may evolve into pure API actions during the later stage of exploration. | format_text_in_word | combine select_text and a series of UI actions related to text styling. |
B.3 Initial Skill Repository
As the foundation for the interaction between the agent and the environment, before exploring and mining skills, we pre-defined some basic actions as the initial skill library shown in Table 7. These actions are derived from UFO Zhang et al. (2024a).
| Name | Description | Example |
| set_edit_text | The function to Set the edit text of the control element, can use to input content on Edit type controls. | set_edit_text(executor, args_dict="control_id":’119’, "control_name":"Edit", ’text’:"hi there") |
| select_text | A function to select the text with the specified text content. | select_text(executor, args_dict="text":"hello") |
| select_table | A function to select the table with the specified number. | select_table(executor, args_dict="number":1) |
| type_keys | A function to Type in keys on control item.Used to enter shortcuts and so on. | type_keys(executor, args_dict="control_id":’119’, "control_name":"Edit", "text": "VK_CONTROL down", "newline": False) |
| click_input | A function to Click the control element.Usually be used to switch to different ribbon,click the buttons in menu. | click_input(executor, args_dict="control_id":"12", "control_name":"Border", ’button’:"left",’double’:False) |
| wheel_mouse_input | A function for Wheel mouse input on the control element. | wheel_mouse_input(executor, args_dict="control_id":"12", ’wheel_dist’:-20) |
B.4 Skill Translation Example
The following is a specific example regarding the generation and translation of skills in Table LABEL:tab:skill-trans.
| Generated Skill by Generator Agent | Translated Skill by Translator Agent |
| ⬇ code: |- from select_text import select_text def hightlight_text(executor, args_dict: dict): """ Highlight the text. :param executor: The executor object. :param args_dict: The arguments of the highlight text method in a dict - text: the text to highlight - home_control_id: the id of the Home ribbon. :return: The result of the action.True or False. """ try: select_text(executor, args_dict) home_control_id = args_dict.get("home_control_id", -1) home_ribbon_control = executor.get_target_control_by_args( control_args={"control_id":home_control_id}) executor.atomic_execution( control=home_ribbon_control, method_name="click_input", args={ "button": "left", "double": False } ) hightlight_button = executor.get_target_control_by_uuid( "uuid_control.json", "96d5a516-35af-459b-b1c4-34aced6acb0b") executor.atomic_execution( control=hightlight_button, method_name="click_input", args={ "button": "left", "double": False } ) except Exception as e: print(f"error:{str(e)}") return False return True description: |- A function to highlight the text. Note: the args_dict should include "home_control_id" field of "Home" ribbon control id and "text" field of the target text to highlight. example: |- highlight_text(executor, args_dict={"home_control_id": 27, "text": "Hello"}) | ⬇ code: |- from select_text import select_text def highlight_text(executor, args_dict: dict): """ Highlight the text. :param executor: The executor object. :param args_dict: The arguments of the highlight text method in a dict - text: the text to highlight :return: The result of the action. True or False. Example: >>> highlight_text(executor, args_dict={"text": "hello"}) """ try: select_text(executor, args_dict) executor.app.Selection.Range.HighlightColorIndex = 7 # 7 represents yellow in Word except Exception as e: print(f"error: {str(e)}") return False return True description: |- The function is to highlight the text in the word document. It first finds the text to be highlighted, then selects the text, and finally highlights the text with yellow color. example: |- highlight_text(executor, args_dict={"text": "hello"}) |
B.5 Explored Skill Examples
Here are some explored skill examples of different hierarchies shown in Table LABEL:tab:explored.
| Hierarchy | Name | Description | Example |
| 1 | activate_dictation | The function is to activate dictation in Microsoft Word. It is equal to the Dictate button in the Voice group to start dictation. | activate_dictation(executor) |
| 2 | align_text | The function aligns the text in a Microsoft Word document. It first selects the text, then applies the desired alignment (left, center, right, justify) using the Word API. | align_text(executor, args_dict="text": "hello", "alignment": "center") |
| 3 | apply_text_style | A function to edit a text with specified text, font size, font name. The title is set in the center. | apply_text_style(executor, args_dict="text":"Hello", "font_name":"Arial", "font_size":13) |
B.6 Skill Dynamic Validator
Dynamic validation evaluates the performance of a skill in actual tasks and environments through two specialized agents as detailed below:
Validator Agent: When a skill is submitted for validation, this agent analyzes the skill’s code, usage examples, and description to generate appropriate test tasks. Create test parameters consistent with the skill’s usage examples and execute the skill in randomly generated Word documents to assess its functionality.
Evaluator Agent: After the Validator Agent executes the target skill, the Evaluator Agent examines the execution logs, document modifications, and final state of the Word document. Evaluate whether the task was successfully completed according to the intended functionality. Only skills that pass this evaluation are eventually added to the skill library.
Appendix C User Study
C.1 Experiment Procedure
The entire user experiment lasted for 30 minutes. During the preparation phase, we sampled five different tasks in Microsoft Word from both official Word documentation and GPT-generated results. Those tasks were categorized into two levels of difficulty: low difficulty (L1) and high difficulty (L2), based on factors such as the number of UI interactions required, the depth of the UI functions, and the number of ribbon switches. Our experimental results also confirmed that tasks in L2 are indeed more difficult than tasks in L1. In the subsequent discussion, we will simply refers tasks in different categories as L1 tasks and L2 tasks. Additionally, we designed a user information form to collect participants’ background information, including their familiarity with Microsoft Word.
We provided users with a simple web interface during the formal experiment, which consisted of two stages. In Stage 1, participants received a pre-printed task list including both L1 and L2 tasks. Based on the task ID displayed on the webpage, participants were asked to read the task requirements, click the "start" button, complete the task in the automatically opened Word document, and click "Finish." upon task completion. In Stage 2, participants were instructed to use both the UI Agent and AXIS to assist them in completing the Word tasks. The corresponding webpage were featured with both input fields and buttons for activating the two Agents. The participants need to enter task description to command the Agents to complete the tasks. Throughout the formal experiment, all task execution processes were recorded for subsequent analysis. After completing all assigned tasks manually or with the assistant of agents, four different post-task questionnaires were displayed on the experimental webpage to survey users’ subjective experiences.
C.2 Participants Recruitment
We recruited candidates by posting on social media. 20 individuals were randomly selected as participants for the experiment from the list of candidates who confirmed their willingness to participate. Our participants ranged in age from 18 to 40 years with educational backgrounds spanning from undergraduate to postgraduate levels. Their occupations included engineers, students, researchers, and full-time homemakers, among others. 100% of the participants had some experience with Microsoft Word with varying levels of proficiency and different usage frequency ranging from daily to monthly. The user experiment lasted 30 minutes on average per participant and each participant received 50 CNY as compensation.
C.3 User Study Web Interface
During the user study, we provided participants with a web interface to control the user study procedure. Below are some screenshots of the web interface.
C.4 User Study Tasks
We sampled five tasks about Microsoft Word in user study which were classified into low difficulty (L1) and high difficulty (L2) following the same criteria as tasks in feasibility study. Here are the detailed tasks in Table LABEL:tab:exp-tasks.
| Task id | Task description | Difficulty level |
| 1 | Here is an article, type in a title "Impossible Friendship between mouse and cats" and set the title in the center with "Arial" type of 20 font size. | L1 |
| 2 | Insert a header named "header" and a footer named "footer". | L1 |
| 3 | Change the titles style of each sections into heading1 style. | L1 |
| 4 | I want to make a special format for company: insert a 2x2 table, then change the paper size in Word to A4, change the text direction to vertical and add water mark with confidential 1 type. | L2 |
| 5 | Insert 2 shapes into document:(1) Insert a rectangle with a width and height of 1 inch, and set the fill color to red. (2) Insert a circle with a width and height of 1 inch, and set the fill color to yellow. | L2 |
C.5 User Study Survey Form
To obtain subjective metrics and analyze the results to address our research questions, we included several questionnaires in the user study which are listed below:
C.5.1 Cognitive load related forms
The cognitive load-related forms include the NASA-TLX survey and the learning effort survey, which participants filled out after completing tasks in both manual mode and agent mode.
C.5.2 Human Preference related forms
The forms related to human preferences include surveys on perceived speed, fluency, reliability, decision consistency, and UI dependency.
Appendix D Feasibility Study
In the feasibility study, we randomly sampled 50 tasks from the WikiHow page ’Use Microsoft Word’ and the official Microsoft Word website. To increase the task difficulty, some of the 50 tasks were composed of smaller sub-tasks, thus increasing the number of steps required for completion. As the tasks sampled in user study, the 50 tasks were also divided into 2 levels of difficulty: low difficulty (L1) and high difficulty (L2), based on three key factors for task difficulty classification:
Number of UI Interactions Required: Measured by the procedural steps (validated against manuals or user testing), shown quantitatively in Table LABEL:tab:50tasks-steps.
Depth of UI Functions: Evaluated via the Control Hierarchy Tree (rooted at the Ribbon level, depth = 0). For example, the task Apply text highlighting (Home → Text Highlight Color) has a depth = 1.
Ribbon Switches Counts: Refers to cross-tab interactions.
A task is classified as L2 difficulty only if it satisfies all three criteria simultaneously: UI interactions, Maximum control depth and Ribbon switch.
Table LABEL:tab:50tasks-steps has shown the distribution of the required execution steps (i.e., the number of steps a human would typically need to perform through UI) of the 50 tasks, along with the number of tasks in different difficulty levels.
| Steps | Tasks Number | Difficulty Level |
| 1 | 3 | L1: 3 L2: 0 |
| 2 | 9 | L1: 9 L2: 0 |
| 3 | 23 | L1: 14 L2: 9 |
| 4 | 12 | L1: 0 L2: 12 |
| 5 | 1 | L1: 0 L2: 1 |
| 8 | 1 | L1: 0 L2: 1 |
| 10 | 1 | L1: 0 L2: 1 |
Appendix E Extensive Applications of AXIS
E.1 AXIS help to digest unnecessary Application UIs
To build new APIs on top of existing API and UI functions, AXIS leverages a LLM-powered self-exploration framework to identify all control elements within an application that can be converted into APIs. This exploration procedure helps uncover potentially unnecessary UI elements or redundant UI designs for improvement under the HACI paradigm.
To illustrate this process, in Figure LABEL:fig:lessui, the UI hierarchical relationships between UIs are represented as a tree, in which each node represents a UI element with higher-level UI elements as parent nodes and lower-level ones as child nodes. We further use red nodes to represent UI locations that can be API - ified after explored by AXIS, and use blue nodes to represent general UI elements. Red nodes imply that the UI controls at these locations can be replaced by APIs. These red - marked UI elements can be described using natural language. In contrast, blue nodes indicate that the corresponding controls are either difficult to describe in language or lack corresponding triggering APIs. In this example, the root node that represents the "Home" tab is a blue node as not all its sub-UI nodes are red (API-ified). However, the second-level node "Highlight Color" (node 2-2) and all its third-level child nodes can all be API-ified and are colored in red. Generally, we define a node N as non-essential if this node along with all their child nodes can all be API-ified:
Unlike the HCI paradigm that emphasizes the interactions between human and interfaces, in the future Agent OS powered by LLM-based agents, non-essential UI elements can be simplified or even eliminated from the application interface, with their original functions replaced by the API calls. By categorizing UI elements as essential or non-essential, AXIS can provide valuable insights on how the UI might be improved and re-designed in an agent-based system for the application providers.
E.2 Turn An Application into an Agent
In the experiment section, we use Microsoft Word to illustrate how to explore and construct new API agents using the AXIS framework. It is worthy noting that the AXIS framework is highly adaptable and scalable, and can be extended to any new application with a basic API and documentation support. Specifically, to adapt AXIS, the application provider needs to supplement operational manuals on the applications as well as the following interfaces:
Environment State Interface for obtaining information about the state of the environment.
Basic Action Interface for supporting basic interactions with the environment.
Starting from those basic resources, AXIS can automatically and continuously explore the applications, discover new skills, and extend its functionalities. This adaptability also means that AXIS can be integrated into various software environments to enhance functionality and user experience with API-driven interactions.